√log 10 x +log 10√ x = - ½

√log10x +log10√x = - ½

Grade:11

1 Answers

Nehal Wani
14 years ago

This equation can be written as (log10x)1/2+log10x1/2=-1/2

or

(log10x)1/2+1/2log10x=-1/2

Put (log10x)=t2

Now

t+t2/2=-1/2

(t+1)2=0

t=-1

(log10x)=1

x=101=10